E6-17 (Static) Analyzing Break-Even and Target Profit [LO 6-1, 6-2]

Tommy’s Tile Service is planning on purchasing new tile cleaning equipment that will improve their ability to remove tough stains from ceramic tiles. The company’s contribution margin is 30% and its current break-even point is $250,000 in sales revenue. Purchasing the new equipment will increase fixed costs by $7,500.

Required:

1. Determine the company’s current fixed costs.

2. Determine the company’s new break-even point in sales.

3. After the purchase of the equipment, how much revenue does the company need to generate a profit of $100,000?
Answer #1

Answers :

1) 75,000

2) 275,000

3) 608,333.33

1) Current Fixed cost = Break even point * CM Ratio = 250,000*0.30 = $75,000(Answer )

2) New Break Even Point = ( 75000+7500)/0.30 = $275,000(Answer)

3) REvenue to generate the Profit of $ 100,000= (Fixed cost + Profit ) / CM Ratio

= (75000+7500+100000)/0.30 = $608,333,33(Answer)